# On the optimality of coin-betting for mean estimation

### v2

@@ 2 Algorithmic mean estimation
Fix a Borel set $\mathcal{X}\subseteq\mathbb{R}$. Let $(X_{t})_{t\geq 1}$ be a sequence of independent random variables drawn from an unknown fixed distribution $P^{\star}\in\mathcal{P}_{\mathcal{X}}$, whose mean $\mu^{\star}\in(0,1)$ has to be estimated. Let $\mathcal{F}=(\mathcal{F}_{t})_{t\geq 0}$ represent the natural filtration generated by this sequence, where $\mathcal{F}_{t}=\sigma(X_{1},\dots,X_{t})$ captures all information available up to time $t$. Fix a confidence level parameter $\delta\in(0,1)$. A confidence sequence $(S_{t})_{t\geq 1}$ is a sequence of random sets such that the sequence of events ($(\{\mu^{\star}\in S_{t}\})_{t\geq 1}$ is adapted with respect to the filtration $\mathcal{F}$ (i.e., for all $t\geq 1$, $\{\mu^{\star}\in S_{t}\}$ is $\mathcal{F}_{t}$-measurable) and satisfies the property
@@
We stress here that, for the time being, we assume that the sequence $(X_{t})_{t\geq 1}$ is i.i.d., and we shall not challenge this assumption, regardless of whether or not the observed data appear to exhibit such behaviour. In a way, this perspective aligns naturally with the goal of constructing confidence sequences, as simultaneously testing for the mean and the i.i.d. assumption might lead to rejecting everything and returning empty confidence sets because “the data do not look i.i.d.”. Under the i.i.d. model, mean estimation is well posed: we assume the existence of a fixed distribution $P^{\star}$ independently generating each observation, and our task is to estimate its mean. We remark that the independence assumption could be relaxed by only asking that each $X_{t}$ has fixed conditional mean $\mu^{\star}$ (to be estimated) given the past $\mathcal{F}_{t-1}$. The arguments we present next for the i.i.d. setting carry over with essentially no change to this less restrictive case. However, to keep the exposition clearer, we focus here solely on the independent case. The conditional setting will be addressed in the second part of this work, where we consider the more general scenario in which dependencies between observations are not assumed but are instead challenged by the statistical test.
2.1 Coin-betting confidence sequence
Recently, Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023) obtained some of the tightest known confidence sequences for bounded random variables via an algorithmic approach. The idea behind both papers is to set up a sequential coin-betting game for each candidate value $\mu\in(0,1)$ for the mean, and to exclude from the confidence sets those values $\mu$ that allow the player to earn unlikely high rewards.
Definition 1 (Coin-betting game).
Fix $\mu\in(0,1)$ and let $I_{\mu}=[(\mu-1)^{-1},\mu^{-1}]$. Consider the following sequential procedure. At each round $t\geq 1$, a player
  • picks [^2] We implicitly assume a measurable selection of $\beta_{t}$, namely the mapping $(x_{1},\dots x_{t-1})\mapsto\beta_{t}$ is Borel measurable. $\beta_{t}\in I_{\mu}$, based solely on the past observations $x_{1},\dots,x_{t-1}$;
  • observes a new data-point $x_{t}\in[0,1]$;
  • gets a reward $\log\big{(}1+\beta_{t}(x_{t}-\mu)\big{)}$.
Now, for each $\mu\in(0,1)$, consider a coin-betting game where at each round the player observes a data-point from the sequence $(X_{t})_{t\geq 1}$ drawn from $P^{\star}$. Let $R_{n}(\mu)=\sum_{t=1}^{n}\log(1+\beta_{t}(X_{t}-\mu))$ denote the cumulative reward at round $n$. Then, via standard martingale concentration arguments (which will become clear later) one can show that the sets $S_{n}=\big{\{}\mu\in(0,1)\,:\,R_{n}(\mu)\leq\log\tfrac{1}{\delta}\big{\}}$ define a valid confidence sequence for $\mu^{\star}$.
2.2 Sequential testing game
The coin-betting approach is a particular instance of constructing confidence sequences via a sequential hypothesis testing procedure based on e-variables. Here, we call hypothesis an arbitrary subset $\mathcal{H}$ of $\mathcal{P}_{[0,1]}$. An e-variable (for $\mathcal{H}$) is any non-negative Borel measurable function $E:[0,1]\to[0,+\infty)$, such that
$\mathbb{E}_{P}[E]\leq 1\,,\qquad\forall P\in\mathcal{H}\,.$
We denote as $\mathcal{E}_{\mathcal{H}}$ the set of all the e-variables with respect to $\mathcal{H}$, and we call e-class any subset of $\mathcal{E}_{\mathcal{H}}$. We remark that $\mathcal{E}_{\mathcal{H}}$ is never empty, since the constant function one is always an e-variable, independently of $\mathcal{H}$.
Definition 2 (Testing-by-betting game).
Fix a hypothesis $\mathcal{H}\subseteq\mathcal{P}_{[0,1]}$ and a non-empty e-class $\mathcal{E}\subseteq\mathcal{E}_{\mathcal{H}}$. We call $\mathcal{E}$-restricted testing-by-betting game (on $\mathcal{H}$) the following sequential procedure. At each round $t\geq 1$,
  • the player picks [^3] Again, we assume that $E_{t}$ is picked in a measurable fashion, namely the mapping $(x_{1},\dots x_{t})\mapsto E_{t}(x_{t})$ is Borel measurable. an e-variable $E_{t}\in\mathcal{E}$ based solely on the past observations $x_{1},\dots,x_{t-1}$;
  • the player observes a new data-point $x_{t}\in[0,1]$;
  • the player earns a reward $\log E_{t}(x_{t})$.
The above game is an instance of e-variable testing (see Ramdas and Wang, 2024, for a recent overview), where one designs a test that rejects the hypothesis $\mathcal{H}$ whenever the total reward earned by the player gets excessively high. This procedure is justified by the fact that, if the data-points observed during the game were independently drawn from $P\in\mathcal{H}$, then the cumulative reward would be unlikely to grow very large.
Lemma 1.
Let $\mathcal{H}\subseteq\mathcal{P}_{[0,1]}$ and consider a sequence $(X_{t}^{\prime})_{t\geq 1}$ of independent draws from some $P\in\mathcal{H}$. Fix $\delta\in(0,1)$ and $\mathcal{E}\subseteq\mathcal{E}_{\mathcal{H}}$. Consider an $\mathcal{E}$-restricted testing-by-betting game, where the observations are the sequence $(X_{t}^{\prime})_{t\geq 1}$. Let $R_{n}=\sum_{t=1}^{n}\log E_{t}(X_{t}^{\prime})$ represent the player’s cumulative reward at round $n$. We have
Proof.
The result follows directly from Ville’s inequality, since $M_{n}=\prod_{t=1}^{n}E_{t}(X_{t}^{\prime})$ defines a non-negative super-martingale with respect to the natural filtration of the process $(X_{t}^{\prime})_{t\geq 1}$, with $M_{0}\equiv 1$. ∎
2.3 Mean estimation via sequential betting
We can leverage the testing-by-betting game of Definition 2 to obtain a confidence sequence for the mean $\mu^{\star}\in(0,1)$ of $P^{\star}$. For $\mu\in(0,1)$, define the hypothesis
$\mathcal{H}_{\mu}=\{P\in\mathcal{P}_{[0,1]}\,:\,\mathbb{E}_{P}[X]=\mu\}\,,$ (1)
@@ 2.3 Mean estimation via sequential betting
which contains all probability measures on $[0,1]$ with mean $\mu$. For each $\mu$, we fix an e-class $\mathcal{E}_{\mu}\subseteq\mathcal{E}_{\mathcal{H}_{\mu}}$ and we consider an $\mathcal{E}_{\mu}$-restricted testing-by-betting game on $\mathcal{H}_{\mu}$, where the player observes $(X_{t})_{t\geq 1}$, the sequence of draws from $P^{\star}$. For each one of these games, denote as $R_{n}(\mu)$ the player’s cumulative reward at round $n$.
Proposition 1.
@@
by Lemma 1, as the observations are drawn from $P^{\star}\in\mathcal{H}_{\mu^{\star}}$. ∎
We remark that the approach to derive confidence sequences via the coin-betting game, proposed by Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023), is a specific instance of the framework that we have just described. Indeed, letting $E_{t}:x\mapsto 1+\beta_{t}(x-\mu)$, it is straightforward to verify that $E_{t}$ is non-negative on $[0,1]$, due to the restriction $\beta_{t}\in I_{\mu}=[(\mu-1)^{-1},\mu^{-1}]$ in Definition 1. Moreover, for any $P\in\mathcal{H}_{\mu}$, we have $\mathbb{E}_{P}[E_{t}]=1$, which implies that $E_{t}$ is an e-variable for $\mathcal{H}_{\mu}$. Finally, the reward in the coin-betting game is precisely equal to $\log E_{t}(x_{t})$. Hence, for the hypothesis $\mathcal{H}_{\mu}$, the coin-betting game matches exactly the testing game of Definition 2, restricted to the coin-betting e-class
Ideally, one aims to construct tight confidence sets. However, achieving this depends on the strategy employed in the testing games. For example, stubbornly playing $E_{t}\equiv 1$ at all rounds in every game would produce valid confidence sets ($S_{n}=(0,1)$, for every $n$) of no practical interest. Indeed, a player’s strategy is most effective when it can rapidly increase the cumulative reward, whenever possible, as this allows to fast exclude redundant points from the confidence sets. In short, the highest the rewards, the tightest the confidence sequence. With this in mind, it is clear that carelessly restricting the player’s choice to a subset $\mathcal{E}_{\mu}$ of $\mathcal{E}_{\mathcal{H}_{\mu}}$ could be highly detrimental, as it might force the player to adopt poor strategies (e.g., consider the trivial case $\mathcal{E}_{\mu}=\{1\}$). This point naturally raises the question: “Does restricting the player to the coin-betting e-class (2) loosen the confidence sequence?”. Interestingly, for the round-wise testing-by-betting framework that we are considering the answer turns out to be negative.
3 Majorising e-classes and optimal e-class
We introduce a partial ordering on the set of real functions on a set $\mathcal{X}$. Given two functions $f,f^{\prime}:\mathcal{X}\to\mathbb{R}$, we say that $f$ majorises $f^{\prime}$, and write $f\succeq f^{\prime}$, if $f(x)\geq f^{\prime}(x)$ for all $x\in\mathcal{X}$. If $f\succeq f^{\prime}$ and there exists a point $x\in\mathcal{X}$ such that $f(x)>f^{\prime}(x)$, we say that $f$ is a strict majoriser of $f^{\prime}$, and we write $f\succ f^{\prime}$.
Definition 3.
$E\in\mathcal{E}_{\mathcal{H}}$ is called a maximal e-variable when there is no $E^{\prime}\in\mathcal{E}_{\mathcal{H}}$ such that $E^{\prime}\succ E$.
Next, we introduce a way to compare different e-classes.
Definition 4.
Given two e-classes $\mathcal{E}$ and $\mathcal{E}^{\prime}$, we say that $\mathcal{E}$ majorises $\mathcal{E}^{\prime}$ if, for any $E^{\prime}\in\mathcal{E}^{\prime}$, there is an e-variable $E\in\mathcal{E}$ such that $E\succeq E^{\prime}$. We call majorising e-class any e-class that majorises $\mathcal{E}_{\mathcal{H}}$.
Lemma 2.
Every majorising e-class contains all the maximal e-variables.
Proof.
Let $E\in\mathcal{E}_{\mathcal{H}}$ be maximal and $\mathcal{E}$ a majorising e-class. There must be $E^{\prime}\in\mathcal{E}$ such that $E^{\prime}\succeq E$, but since $E$ is maximal it has to be that $E=E^{\prime}$. Hence, $E\in\mathcal{E}$. ∎
The significance of the notion of majorising e-class for our problem is straightforward: if $\mathcal{E}$ majorises $\mathcal{E}^{\prime}$, then any strategy in an $\mathcal{E}^{\prime}$-restricted game can be matched or outperformed (in terms of rewards) by a corresponding strategy in the $\mathcal{E}$-restricted game, regardless of the sequence of observations. This allows us to compare how the restriction to different e-classes affects the testing-by-betting game of Definition 2.
Definition 5.
If a majorising e-class is contained in every other majorising e-class, we call it optimal.
For any $\mathcal{H}$, a majorising e-class always exists, as $\mathcal{E}_{\mathcal{H}}$ itself is a majorising e-class. However, an optimal e-class may not exist. [^4] We refer the reader to the follow-up work Clerico (2024) for an example of non-existence of the optimal e-class. Proposition 2 therein shows that =$\mathcal{H}=\{P\in\mathcal{P}_{[0,1]}\,:\,P(\{0\})\geq 1/2\}\cup\{U_{[0,1]}\}$ (with $U_{[0,1]}$ the uniform distribution on $[0,1]$) is a hypothesis for which the set of maximal e-variables is not a majorising e-class. Hence, $\mathcal{H}$ does not admit an optimal e-class. Next, we state a sufficient and necessary condition for the existence of optimal e-classes.
Lemma 3.
Proof.
Denote as $\hat{\mathcal{E}}$ the set of all the maximal e-variables. Now, assume that there exists an optimal e-class $\mathcal{E}$. First, we show that all its elements must be maximal. For $E\in\mathcal{E}$, consider any element $E^{\prime}\in\mathcal{E}_{\mathcal{H}}$ such that $E^{\prime}\succeq E$. We can construct an e-class $\mathcal{E}^{\prime}$ replacing $E$ with $E^{\prime}$ in $\mathcal{E}$, namely =($\mathcal{E}^{\prime}=(\mathcal{E}\setminus\{E\})\cup\{E^{\prime}\}$. Since $E^{\prime}\succeq E$, it is clear that $\mathcal{E}^{\prime}$ majorises $\mathcal{E}$. It follows that $\mathcal{E}^{\prime}$ is a majorising e-class, as $\mathcal{E}$ is. Since $\mathcal{E}$ is optimal, $\mathcal{E}\subseteq\mathcal{E}^{\prime}$, which implies $E=E^{\prime}$. In particular, $E$ does not have any strict majoriser, and so it is maximal. Hence, $\mathcal{E}\subseteq\hat{\mathcal{E}}$. As $\mathcal{E}\supseteq\hat{\mathcal{E}}$ by Lemma 2, we conclude that $\mathcal{E}=\hat{\mathcal{E}}$, and so $\hat{\mathcal{E}}$ is a majorising e-class.
Now, the final statements are a trivial consequence of what was shown above and Lemma 2. ∎
Let us emphasise once more that, from our discussion thus far, it is clear that restricting the testing-by-betting game of Definition 2 to a majorising e-class does not hinder the performance of the player, as for any unrestricted strategy $(E_{t})_{t\geq 1}$ they can always pick a restricted strategy $(E_{t}^{\prime})_{t\geq 1}$, whose cumulative rewards inevitably match or outperform those of $(E_{t})_{t\geq 1}$, regardless of the sequence of observations. From a practical perspective, identifying the optimal e-class, when it exists, greatly simplifies the design of an effective strategy by narrowing the player’s choice to the best possible e-variables. Specifically, if the optimal e-class exists and a player chooses an e-variable $E_{t}$ outside of it, they could always have picked an alternative $E_{t}^{\prime}\succ E_{t}$, within the optimal e-class, whose reward is never worse than that of $E_{t}$ and is strictly higher for at least one possible value that $x_{t}$ might take. Conversely, when a player selects an e-variable from the optimal e-class, no other choice can be guaranteed to be better before observing $x_{t}$, since the player’s pick is a maximal e-variable. As a straightforward consequence, within this round-wise testing-by-betting approach, the optimal approach to obtain confidence sequences consists in restricting game to the optimal e-class, for each mean candidate value $\mu$. We show next that this coincides precisely with the coin-betting formulation.
4 Optimality of the coin-betting e-class
For any $\mu\in(0,1)$, define the hypothesis $\mathcal{H}_{\mu}$ as in (1). We now show that the optimal e-class for $\mathcal{H}_{\mu}$ exists and coincides with the coin-betting e-class $\mathcal{E}^{\mathrm{cb}}_{\mu}$, defined in (2).
Lemma 4.
Fix $\mu\in(0,1)$ and consider the function $F_{\mu}:[0,1]\to[1,+\infty)$ defined as
[…]
For any $x\in[0,1]$ there is $P_{x}\in\mathcal{H}_{\mu}$ such that $P_{x}(\{x\})=1/F_{\mu}(x)>0$. Moreover, $F_{\mu}\succeq E$ for every $E\in\mathcal{E}_{\mathcal{H}_{\mu}}$.
Proof.
For $x\geq\mu$, let $P_{x}=\frac{\mu}{x}\delta_{x}+(1-\frac{\mu}{x})\delta_{0}$. Then, $P_{x}\in\mathcal{H}_{\mu}$ and $P_{x}(\{x\})=\mu/x=1/F_{\mu}(x)$. With a similar reasoning, if $x<\mu$ we can find a measure $P_{x}$ in $\mathcal{H}_{\mu}$, supported on $\{x,1\}$, with mass $1/F_{\mu}(x)$ on $x$.
Now, to check that $F_{\mu}$ majorises all the e-variables, fix $E\in\mathcal{E}_{\mathcal{H}_{\mu}}$ and $x\in[0,1]$. Consider a $P_{x}\in\mathcal{H}_{\mu}$ with mass $1/F_{\mu}(x)$ on $x$. Then, $1\geq\mathbb{E}_{P_{x}}[E]\geq P_{x}(x)E(x)=E(x)/F_{\mu}(x)$, and we conclude. ∎
Theorem 1.
Proof.
$B_{0}=\left\{\beta\in I_{\mu}\;:\;\inf_{x\in[0,\mu)}\big{(}E_{\beta}(x)-E(x)\big{)}\geq 0\right\}\quad\text{and}\quad B_{1}=\left\{\beta\in I_{\mu}\;:\;\inf_{x\in(\mu,1]}\big{(}E_{\beta}(x)-E(x)\big{)}\geq 0\right\}\,,$
where we recall that $I_{\mu}=[(\mu-1)^{-1},\mu^{-1}]$ and $E_{\beta}:x\mapsto 1+\beta(x-\mu)$. Both sets are closed and convex (as they are intersections of closed and convex sets). By Lemma 4, $(\mu-1)^{-1}\in B_{0}$ and $\mu^{-1}\in B_{1}$, so $B_{0}=[(\mu-1)^{-1},\beta_{0}]$ and $B_{1}=[\beta_{1},\mu^{-1}]$, for some $\beta_{0}$ and $\beta_{1}$ in $I_{\mu}$. We will now show that $B_{0}\cap B_{1}\neq\varnothing$, or equivalently that $\beta_{0}\geq\beta_{1}$. Assume that this was not the case and $\beta_{0}<\beta_{1}$. Let $\beta^{\star}\in(\beta_{0},\beta_{1})$. Then, $\beta^{\star}\notin B_{0}$ and $\beta^{\star}\notin B_{1}$. In particular, there are $u_{0}\in[0,\mu)$ and $u_{1}\in(\mu,1]$ such that $E(u_{0})>E_{\beta^{\star}}(u_{0})$ and $E(u_{1})>E_{\beta^{\star}}(u_{1})$. As $\mu\in(u_{0},u_{1})$, there is $\hat{P}\in\mathcal{H}_{\mu}$ with support $\{u_{0},u_{1}\}$. Note that $\mathbb{E}_{\hat{P}}[E_{\beta^{\star}}]=1+\beta^{\star}(\mathbb{E}_{\hat{P}}[X]-\mu)=1$. But $E$ is strictly larger than $E_{\beta^{\star}}$ on $\mathrm{Supp}(\hat{P})$, and so $\mathbb{E}_{\hat{P}}[E]>1$, which is a contradiction since $E$ is an e-variable. We conclude that $\beta_{0}\geq\beta_{1}$ and so there exists $\hat{\beta}\in B_{0}\cap B_{1}$. By construction, $E_{\hat{\beta}}\in\mathcal{E}^{\mathrm{cb}}_{\mu}$ and $E_{\hat{\beta}}(x)\geq E(x)$ for all $x\in[0,1]$ different from $\mu$. For $x=\mu$, by Lemma 4 we have $E(\mu)\leq F_{\mu}(\mu)=1=E_{\hat{\beta}}(\mu)$, and so $E\preceq E_{\hat{\beta}}$. As the choice of $E$ was arbitrary, $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class.
Now that we have established that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class, by Lemma 3 we only need to show that all its elements are maximal to conclude that it is optimal. Fix $E\in\mathcal{E}^{\mathrm{cb}}_{\mu}$, and consider an e-variable $\hat{E}\in\mathcal{E}_{\mathcal{H}_{\mu}}$ such that $\hat{E}\succeq E$. Fix any $x\in[0,1]$. By Lemma 4, there is $P\in\mathcal{H}_{\mu}$ such that $P(\{x\})>0$. Since $E\in\mathcal{E}^{\mathrm{cb}}_{\mu}$, we have $\langle P,E\rangle=1$, and so $0\leq P(\{x\})(\hat{E}(x)-E(x))\leq\mathbb{E}_{P}[\hat{E}-E]=\mathbb{E}_{P}[\hat{E}]-1\leq 0$. Since $P(\{x\})>0$, we get $\hat{E}(x)=E(x)$ and, $x$ being arbitrary, $\hat{E}=E$. Hence $E$ is maximal, as it has no strict majoriser. ∎

### v4

@@ 5 From mean testing to confidence sequences
As usual, $(X_{t})_{t\geq 1}$ is a sequence of independent draws from $P^{\star}\in\mathcal{P}(\mathcal{X})$, whose mean $\mu^{\star}\in(0,1)$ has to be estimated. Let $\mathcal{F}=(\mathcal{F}_{t})_{t\geq 0}$ represent the natural filtration generated by $(X_{t})_{t\geq 1}$, where $\mathcal{F}_{t}=\sigma(X_{1},\dots,X_{t})$ captures all information available up to time $t$. Fix a confidence level parameter $\delta\in(0,1)$. A confidence sequence $(S_{t})_{t\geq 1}$ is a sequence of random sets [^6] Here, by random sets we simply mean sets that depend on the sequence of random observations $(X_{t})_{t\geq 1}$. such that the sequence of events ($(\{\mu^{\star}\in S_{t}\})_{t\geq 1}$ is adapted to the filtration $\mathcal{F}$ (i.e., for all $t\geq 1$, $\{\mu^{\star}\in S_{t}\}$ is $\mathcal{F}_{t}$-measurable) and satisfies
@@
We can leverage the testing-by-betting game of Definition 1 to obtain a confidence sequence for the mean $\mu^{\star}$ of $P^{\star}$. For $\mu\in(0,1)$, we define the hypothesis $\mathcal{H}_{\mu}$ as in (1), namely $\mathcal{H}_{\mu}$ contains all probability measures on $\mathcal{X}$ with mean $\mu$. For each $\mu$, we fix an e-class $\mathcal{E}_{\mu}\subseteq\mathcal{E}_{\mathcal{H}_{\mu}}$ and we consider an $\mathcal{E}_{\mu}$-restricted testing-by-betting game on $\mathcal{H}_{\mu}$, where the player observes $(X_{t})_{t\geq 1}$, the sequence of draws from $P^{\star}$. For each one of these games, denote as $R_{n}(\mu)$ the player’s cumulative reward at round $n$. We can then construct a confidence sequence as follows.
Proposition 3
@@
by Proposition 1, as the observations are drawn from $P^{\star}\in\mathcal{H}_{\mu^{\star}}$. $\qed$
It is worth stressing that the strength of the resulting confidence sequence depends directly on the power of the underlying sequential tests for each hypothesis $\mathcal{H}_{\mu}$: more powerful tests yield tighter confidence sets. In particular, the optimality result for the coin-betting e-class established in Section 4 implies that, for each $\mu$, the testing game should be restricted to the class $\mathcal{E}^{\mathrm{cb}}_{\mu}$. With this choice, we recover the coin-betting framework to derive confidence sequences leveraged by Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023).
Both Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023) propose explicit strategies for placing bets in the coin-betting game, leading to concrete confidence sequences. While the present work’s focus is on defining the optimal betting game rather than designing specific strategies, it is still useful to look more closely at the approach of Orabona and Jun (2023) to illustrate how the optimality of the coin-betting e-class simplifies the construction of a strategy for the game of Definition 1. They employ a coin-betting strategy based on the universal portfolio algorithm, a special instance of online learning Bayesian aggregation techniques (see, e.g., Chapters 9 and 10 of Cesa-Bianchi and Lugosi, 2006). In this approach, a prior distribution $\rho_{1}$ is fixed over the decision space $I_{\mu}$, and then updated sequentially using observed data. Specifically, at round $t$, one defines
$\mathrm{d}\rho_{t}(\lambda)=\zeta_{t}^{-1}\prod_{i=1}^{t-1}(1+\lambda x_{i})\mathrm{d}\rho_{1}(\lambda)\,,$
where $\zeta_{t}$ is the normalising constant ensuring that $\rho_{t}$ is a probability measure on $I_{\mu}$. The bet $\lambda_{t}$ is then chosen as the posterior mean: $\lambda_{t}=\int_{I_{\mu}}\lambda\mathrm{d}\rho_{t}(\lambda)$.
This strategy can be interpreted as performing Bayesian averaging over the coin-betting e-class $\mathcal{E}^{\mathrm{cb}}_{\mu}$, leveraging its simple and low-dimensional parametric structure. This is a concrete example of how knowing explicitly the optimal e-class can help designing powerful testing-by-betting strategies. Indeed, such approach would not be feasible on the full class of all e-variables, an infinite dimensional functional space where even the definition of a prior can become problematic, if we did not know that the prior should be supported on the coin-betting e-class. Importantly, the optimality of $\mathcal{E}^{\mathrm{cb}}_{\mu}$ ensures that no statistical power is sacrificed by restricting to this class. Notably, a similarly structured averaging strategy over a suboptimal e-class, such as the Hoeffding e-class or its convex hull, would remain well defined and computationally tractable, but lead to strictly worse performance, as shown by Proposition 2.